Suppose that firms A, B, C and D are Bertrand duopolists in the salt industry. The market demand curve can be specified as Q=100-3p, Q=qA+qB+qC+qD.

(The firms choose prices simultaneously.)

The cost to firm A is C(qA)=7qA.

The cost to firm B is C(qB)=3qB.

The cost to firm C is C(qC)=7qC

The cost to firm C is C(qD)=3qD
Answer #1

1a) Form A will earn - (J.A =0) : Ans is $0 1b) Firm B will earn - (J. B =$246 or more) . Ans. is $246 or more (C) Firm c wil

Clarification: --

As Firms A, B, C, and D are Bertrand duopolists and in Bertrand competition, firms produce where P = MC as all the three firms A, C and D have irrelevant cost = $7 while Firm B's MC = $3

  • So simply Firm B will convey the whole sum while the different firms don't make anything.

By the Firm B will pick a worth which is just not actually the Marginal cost of the different firms. Acknowledge it pick an expense of $6.(it could pick $6.9 anyway here I am taking $6 to make it whole number)

eso it Produce e = 100 - 3*6 = 82 . P = $6 TC = 39B AC - TC/98=3 so probruit = (P-ATC)* a = (6-3)* 82 = $246 (birm B will ear